Brooke is 3 year younger than Samantha. In five years the sum of their ages will be 29. Find Brooke's age.
Answer by trevtoes(3)

Brooke = X-3
Samantha =X
(X+5)+(X+5-3)=29
2X+10-3=29
2X+7=29
-7 -7
2x=22
divide both sides by 2
X=11
Brookes age is X-3 so 11-3= 8 years old